### v2

@@ Abstract
Foundation models pretrained on large-scale datasets are revolutionizing the field of computational pathology (CPath). The generalization ability of foundation models is crucial for the success in various downstream clinical tasks. However, current foundation models have only been evaluated on a limited type and number of tasks, leaving their generalization ability and overall performance unclear. To address this gap, we established a most comprehensive benchmark to evaluate the performance of off-the-shelf foundation models across six distinct clinical task types, encompassing a total of 39 specific tasks. Our findings reveal that existing foundation models excel at certain task types but struggle to effectively handle the full breadth of clinical tasks. To improve the generalization of pathology foundation models, we propose a unified knowledge distillation framework consisting of both expert and self-knowledge distillation, where the former allows the model to learn from the knowledge of multiple expert models, while the latter leverages self-distillation to enable image representation learning via local-global alignment. Based on this framework, a Generalizable Pathology Foundation Model (GPFM) is pretrained on a large-scale dataset consisting of 190 million images from around 86,000 public H&E whole slides across 34 major tissue types. Evaluated on the established benchmark, GPFM achieves an impressive average rank of 1.36, with 29 tasks ranked 1st, while the second-best model, UNI, attains an average rank of 2.96, with only 4 tasks ranked 1st. The superior generalization of GPFM demonstrates its exceptional modeling capabilities across a wide range of clinical tasks, positioning it as a new cornerstone for feature representation in CPath.
@@
In recent years, there has been a notable progress in the fields of computer vision and natural language processing driven by self-supervised learning on large-scale datasets. These pretrained models, commonly referred to as foundation models, have garnered significant attention and have exhibited remarkable success across various tasks [28, 29, 30]. In the field of CPath, some efforts [31, 32, 33, 34, 35, 36, 37] have been dedicated to pretraining foundation models that can learn inherent representations of histopathology images, catering to the diverse array of tasks encountered in clinical pathology practice. However, the current foundation models have only been evaluated on a limited type of tasks as shown in Figure 2.a, leaving their overall performance unclear. To comprehensively evaluate these models, we built a most comprehensive benchmark spanning six major clinical task categories, comprising 39 specific tasks. Our findings revealed that the generalization ability of these models is still limited and no single model can effectively address all the tasks as shown in Figure 1.c. It can be seen that UNI [33] achieves the best performance in image retrieval and patch-level (ROI) tissue classification tasks, Phikon [32] performs best in survival analysis and report generation tasks, and CONCH [35] obtains highest performance in WSI classification and visual question answering (VQA) tasks. This can be attributed to the fact that each foundation model is trained using distinct datasets and pretraining strategies, leading to specific advantages for each model within particular datasets. These findings highlight the need for further research to develop more generalizable foundation models that can consistently perform well across the diverse type of clinical tasks encountered in CPath. By addressing this challenge, we can unlock the full potential of the foundation model in CPath.
@@
To improve the generalization of pathology foundation model and enhance the overall performance, an intuitive idea is to leverage the specific strengths of existing models by employing knowledge distillation techniques [38, 39]. Accordingly, we proposed a novel self-supervised learning framework with expert and self knowledge distillation to develop a Generalizable Pathology Foundation Model (GPFM). Based on this pretraining method, we trained GPFM using a diverse dataset of approximately 190 million images from 86,104 public WSIs across 34 major tissue types. The overview of GPFM is presented in Figure 1. With the collected diverse tissues and the indirectly using of the images that used to pretrain expert models (e.g., UNI and CONCH), GPFM exhibits outstanding performance across the established benchmarks, achieving an average rank of 1.36 (29 out of 39 tasks), while the second-best performing model, UNI, achieving an average rank of 2.96 (4 out of 39 tasks). These results demonstrate the efficacy of GPFM as a generalizable foundation model in CPath, showcasing its potential to significantly advance the field. The ability of GPFM to consistently perform well across a diverse type of clinical tasks highlights the benefits of leveraging knowledge distillation to combine the strengths of expert models, ultimately leading to more robust and versatile foundation models for supporting clinicians and improving patient care.
@@ 2 Results
We evaluated various foundation models across 39 tasks, encompassing WSI classification (12 tasks), survival analysis (12 tasks), patch-level (ROI) tissue classification (12 tasks), pathological visual question answering, report generation, and pathological image retrieval. The overall results are presented in Figure 2. Since the tasks involved different types of evaluation metrics, we assessed the overall performance of the foundation models using an average ranking approach and reported the critical difference (CD) diagram [40, 41, 42]. The model with the best performance was ranked 1st, while the model with the lowest performance was ranked 7th. Across all tasks, the GPFM model achieved the top average rank score of 1.36 (best performance 29 out of 39), outperforming the second-best model, UNI, which had a ranking score of 2.96 (best performance 4 out of 39). Furthermore, we conducted the Nemenyi statistical test [40] to assess whether the ranking score of GPFM was significantly different from the other foundation models, as shown in Figure 2.c. The results indicate that GPFM had a statistically significant CD compared to the other six models, while UNI, Phikon, CONCH, and Ctranspath did not show a significant CD.
In addition, we also simply averaged the evaluation metrics across all 39 tasks, as shown in Figure 2.b. The GPFM model achieved an average score of 0.859, while the second-best performing model, Phikon, achieved an average score of 0.849. To further assess the significance of the performance difference, we conducted the Wilcoxon signed-rank test [40] between GPFM and the other foundation models. The results indicate that all p-values were less than 0.001, suggesting that GPFM consistently and significantly outperformed the existing foundation models. It is worth noting that the second-best performing model is different depending on the evaluation perspective. From the ranking standpoint, as discussed earlier, the UNI model was the second-best performer with an average ranking score of 2.96. Considering both the ranking perspective and the average metric aspect, the results clearly indicate that GPFM achieves state-of-the-art performance and is much more generalizable compared to the other foundation models. The detailed results can be found in Extended Data Table 1.
@@ 2.1 WSI Classification
WSI classification is pivotal in accurate cancer diagnosis. It aids in categorizing the specific subtype of cancer, which can be significantly improved by utilizing foundation models. Therefore, it is important to evaluate the representation learning capabilities of different foundation models. In this study, we employed ABMIL [43] to assess the WSI classification performance as previous work [33]. We conducted experiments on a total of 12 tasks and the detailed experimental results are presented in Extended Data Table 2-Table 14.
Across 12 WSI classification tasks, GPFM achieved an average ranking score of 1.08, outperforming the second-best performing model, UNI, which achieved an average ranking score of 3.04.
It’s worth noting that the ranking is based on the Area Under the Curve (AUC) metric, where a higher AUC indicates better performance.
The detailed ranking scores of WSI classification tasks are presented in Figure 2.d.
Furthermore, we evaluated the overall performance using average metrics, including AUC, balanced accuracy, and weighted F1 score.
As shown in Figure 3.a, GPFM achieved the best average AUC of 0.956, representing a 1.4% improvement over the second-best performing model, CONCH, with statistical significance P < 0.001 (+1.4%, CONCH, P<0.001).
Similarly, as depicted in Figure 3.b-c, GPFM achieved the best balanced accuracy of 0.833 (+1.9%, UNI, P<0.001) and the best weighted F1 score of 0.834 (+1.9%, UNI, P<0.001).
The detailed performance of the different foundation models is presented in Figure 3.e-l and Extended Data Figure 9.a-d.
We also visualized the ROC curves of GPFM in Extended Data Figure10 for every classification task.
These results across multiple performance metrics clearly illustrated the great potential and strong generalization ability of the GPFM model in the WSI classification tasks.
GPFM improves cancer diagnosis. Cancer diagnosis using WSIs is very common in CPath. Among 10 diagnosis tasks, GPFM achieved the best performance in 8 tasks including breast cancer (BRACS-3 [44], BRACS-7 [44], and BRCA [45], CAMELYON [46, 47]), lung cancer(CPTAC-NSCLC [45]), renal cell carcinoma (RCC [45]), prostate cancer (PANDA [48]), and
ovarian cancer (UBC-OCEAN, an external dataset [49]).
On the fine-grained brain tumor classification task (EBRAINS, an external dataset [50]), GPFM and UNI tied for first place, while on the lung cancer subtyping based on NSCLC dataset [45], GPFM and CONCH tied for first place.
The results illustrate that GPFM consistently performs best and is more robust than other foundation models in cancer diagnosis tasks.
@@
Following the same evaluation strategy as cancer diagnosis, we also investigated the potential of foundation models in 2 gene mutation prediction tasks.
The GPFM achieved the best performance both in LUAD-TP53 (from TCGA) and Glioma IDH 1 (from TCGA GBMLGG) tasks, with AUCs of 0.855 and 0.998, respectively.
Compared to the second-best model, Phikon, the GPFM showed a 1.2% improvement in AUC. These results illustrate that the GPFM is more generalizable than existing foundation models for gene mutation prediction tasks.
Overall, the results from both the cancer diagnosis and gene mutation prediction experiments demonstrate that the GPFM is more generalizable than existing foundation models. Compared to previous foundation models that did not utilize knowledge distillation, the GPFM was able to integrate knowledge from adopted expert models through a unified knowledge distillation mechanism. This mechanism likely allowed the GPFM to learn from a broader set of data and perspectives, which in turn boosted its performance. The ability of the GPFM to leverage knowledge distillation appears to be a key factor contributing to its superior generalizability compared to other foundation models.
@@ 2.2 Survival Analysis
In the context of clinical trials for oncology, survival analysis is commonly employed, with the time to an event, such as death or disease progression, serving as the primary outcome under investigation [51, 52, 53, 54]. Accurate prediction of a patient’s survival risk can enable more targeted and effective treatment strategies. A robust foundation model is essential for improving the precision of survival risk prediction, ultimately leading to better patient outcomes. To evaluate the performance of various foundation models in survival analysis, we conducted experiments on 12 datasets. Following the methodologies of previous works [55, 52, 54], we adopted the Concordance Index (C-Index) as the evaluation metric to compare the performance of different foundation models. The overall comparison of various foundation models is presented in Figure 4, and the detailed experimental results are reported in Extended Data Table 15 - Table 19. The survival curves of GPFM are provided in Extended Data Figure 11 for exploring the detailed distribution of the results.
Across all 12 survival analysis tasks, the GPFM achieved an impressive average ranking score of 1.42, outperforming the competitors in 7 out of the 12 tasks. In comparison, the second-best performing model, Phikon, attained an average ranking score of 2.08, achieving the best performance in 4 out of the 12 tasks. The detailed ranking scores of the various foundation models are presented in the Figure 2.d. Furthermore, when evaluated using the widely recognized C-Index metric, the GPFM once again emerged as the top performer, achieving an average C-Index of 0.726. This result represents a statistically significant improvement of 0.6% over the Phikon model (P<0.001), further solidifying the generalization capability of GPFM for survival analysis tasks. Among these 12 tasks, the 5-fold cross validation was adopted for 11 of them. For the CPTAC-LUAD dataset, it was evaluated directly using the ABMIL model trained on the TCGA-LUAD dataset. The experimental results showed that the C-Index of Phikon dropped from 0.738 to 0.577, a decrease of 0.161. In contrast, GPFM shows better generalization ability with its C-Index dropping from 0.739 to 0.626, a smaller decrease of 0.113. The detailed comparisons between GPFM and other foundation models are presented in Figure 4.
@@
It is interesting to note that while UNI and CONCH are powerful in WSI classification tasks, they do not perform as well in survival analysis tasks. As shown in Figure 2.d, GPFM and Phikon almost achieved the top performance in all 12 tasks. The experimental results in both WSI classification and survival analysis clearly demonstrate the limited generalization capability of the existing foundation models. This limitation is likely due to the data distribution of their training sets and the pretraining methods they employed. While the existing foundation models lack generalization, they perform exceptionally well on specific type of tasks. By combining their strengths, it may be possible to build a more powerful and versatile model. That is exactly what we have done — by proposing the unified distillation framework in this paper to distill the capabilities of the tasks that the existing models excel at into the GPFM, thereby significantly improving its generalization ability.
@@ 2.3 ROI Classification
The performance of WSI classification is influenced by both the feature extractor (foundation model) and the MIL method. In the WSI classification tasks, we use the ABMIL to evaluate whether the features extracted by the foundation models are discriminative. In Region of Interest (ROI) classification tasks, we can directly evaluate the feature representation abilities of the foundation models without using any MIL method. Specifically, we evaluated the feature representation abilities of the foundation models by linear probe, similar to previous work [56]. The detailed experimental results can be found in the Extended Data Table 20 to Table 32.
Across all 12 ROI classification tasks, GPFM achieved the best ranking score of 1.60, while the second-best performing model, UNI, attained a ranking score of 2.67. The detailed ranking scores are presented in Figure 2.d. Furthermore, when considering the conventional metrics, GPFM attained the highest AUC of 0.955 (+0.2%, UNI, P<0.001), the best weighted F1 score of 0.848 (+0.8%, UNI, P<0.001), and the top balanced accuracy of 0.853 (+0.8%, UNI, P<0.001) on average. The comparison of various foundation models is illustrated in Figure 5. Specifically, GPFM demonstrated the best performance in the CRC-MSI, PanCancer-TIL, PCAM, WSSS4LUAD datasets, with AUC values of 0.812 (+0.2%, CONCH, P<0.001), 0.978 (+0.1%, UNI, P<0.001), 0.988 (+0.6%, UNI, P<0.001), 0.998 (+0.1%, Phikon, P<0.001), respectively. Since ROI classification tasks are relatively easy, GPFM also jointly ranked first with other foundation models in the CRC-100K, CCRCC-TCGA-HEL, and PanCancer-TCGA datasets, achieving AUC values of 0.995 (Ctranspath), 0.997 ( Phikon), 0.999 (Phikon), respectively. We also conducted external validation on Chaoyang, UniToPatho, and BreakHis datasets. Across the three external validation datasets, the GPFM achieved the best average rank of 1.5, while the second best performing model, UNI, obtained an average rank of 2.3, illustrating the generalization capability of GPFM.
@@
Furthermore, to assess the robustness of GPFM in handling images with varying resolutions, we visualized the heatmap of attention scores between the [patch] tokens and [CLS] tokens of the ViT transformer. The visualization result is depicted in Figure 9.l. Across four different resolutions (224x224, 448x448, 896x896, and 1344x1344), we observed consistent attention patterns, which illustrates the robustness of GPFM in accommodating varying resolutions.
An examination of the average ranking score in Figure 5.a reveals that the performance of Phikon, CONCH, and UNI perform closely, with ranking score differences less than 1. This suggests that no single existing model dominates in ROI classification tasks. However, GPFM achieves an average ranking score of 1.60, more than one point ahead of the next best model, demonstrating its strength as a generalizable foundation model. The above results also verify the effectiveness of the proposed unified knowledge distillation approach. By leveraging the knowledge of all adopted expert models, the unified knowledge distillation method is able to outperform the individual expert model.

### v3

@@ Abstract
Foundation models pretrained on large-scale datasets are revolutionizing the field of computational pathology (CPath). The generalization ability of foundation models is crucial for the success in various downstream clinical tasks. However, current foundation models have only been evaluated on a limited type and number of tasks, leaving their generalization ability and overall performance unclear. To address this gap, we established a most comprehensive benchmark to evaluate the performance of off-the-shelf foundation models across six distinct clinical task types, encompassing a total of 72 specific tasks, including slide-level classification, survival prediction, ROI-tissue classification, ROI retrieval, visual question answering, and report generation. Our findings reveal that existing foundation models excel at certain task types but struggle to effectively handle the full breadth of clinical tasks. To improve the generalization of pathology foundation models, we propose a unified knowledge distillation framework consisting of both expert and self-knowledge distillation, where the former allows the model to learn from the knowledge of multiple expert models, while the latter leverages self-distillation to enable image representation learning via local-global alignment. Based on this framework, we curated a dataset of 96,000 whole slide images (WSIs) and developed a Generalizable Pathology Foundation Model (GPFM). This advanced model was trained on a substantial dataset comprising 190 million images extracted from approximately 72,000 publicly available slides, encompassing 34 major tissue types. Evaluated on the established benchmark, GPFM achieves an impressive average rank of 1.6, with 42 tasks ranked 1st, while the second-best model, UNI, attains an average rank of 3.7, with only 6 tasks ranked 1st. The superior generalization of GPFM demonstrates its exceptional modeling capabilities across a wide range of clinical tasks, positioning it as a new cornerstone for feature representation in CPath.
@@
In recent years, there has been a notable progress in the fields of computer vision and natural language processing driven by self-supervised learning on large-scale datasets. These pretrained models, commonly referred to as foundation models (FM), have garnered significant attention and have exhibited remarkable success across various tasks [28, 29, 30]. In the field of CPath, some efforts [31, 32, 33, 34, 35, 36, 37] have been dedicated to pretraining FMs that can learn inherent representations of histopathology images, catering to the diverse array of tasks encountered in clinical pathology practice. However, the current FMs have only been evaluated on a limited type of tasks (Fig. 2a), leaving their overall performance unclear. To comprehensively evaluate these models, we built a most comprehensive benchmark spanning six major clinical task categories (Fig. 1d), comprising 72 specific tasks. Our findings revealed that the generalization ability of these models is still limited and no single model can effectively address all the tasks (Fig. 1d). It can be seen that UNI [33] achieves the best performance in WSI classification, image retrieval, survival analysis, and patch-level (ROI) tissue classification tasks, Phikon [32] performs best in report generation tasks, and CONCH [35] obtains highest performance in visual question answering (VQA) tasks. This can be attributed to the fact that each FM is trained using distinct datasets and pretraining strategies, leading to specific advantages for each model within particular datasets. These findings highlight the need for further research to develop more generalizable FMs that can consistently perform well across the diverse types of clinical tasks encountered in CPath. By addressing this challenge, we can unlock the full potential of the FM in CPath.
@@
To improve the generalization of pathology FM and enhance the overall performance, an intuitive idea is to leverage the specific strengths of existing models by employing knowledge distillation techniques [38, 39]. Accordingly, we proposed a novel self-supervised learning framework with expert and self knowledge distillation to develop a Generalizable Pathology Foundation Model (GPFM). Based on the aforementioned pretraining method, we collected a dataset comprising 95,572 slides, encompassing 34 major tissue types, for the purpose of training and evaluating the GPFM. From this collection, we extracted 190 million patches derived from 72,280 slides to facilitate the pre-training (Fig. 1a). With the collected diverse tissues and the indirectly using of the images that used to pretrain expert models (e.g., UNI and CONCH), GPFM exhibits outstanding performance across the established benchmarks (Fig. 1b-c), achieving an average rank of 1.6, while the second-best performing model, UNI, achieves an average rank of 3.7 (Fig. 2c). These results demonstrate the efficacy of GPFM as a generalizable FM in CPath, showcasing its potential to significantly advance the field. The consistent performance of GPFM across a diverse range of clinical tasks underscores the advantages of employing knowledge distillation to integrate the strengths of specialized expert models. This approach facilitates the development of more robust and versatile foundation models (FMs), thereby enhancing their utility in supporting clinical decision-making and advancing patient care outcomes.
@@ 2 Results
We evaluated various FMs across 72 tasks, encompassing 36 WSI classification tasks, 15 survival analysis tasks, 16 patch-level (ROI) tissue classification tasks, 2 pathological visual question answering task, 2 report generation tasks, and 1 pathological image retrieval task (Fig. 2e-g). Since the tasks involved different types of evaluation metrics, we assessed the overall performance of the FMs using an average ranking approach and reported the critical difference (CD) diagram [40, 41, 42]. The model with the best performance was ranked 1st, while the model with the lowest performance was ranked 9th. Across all tasks, the GPFM model achieved the top average rank score of 1.6 (ranked first in 42 tasks), outperforming the second-best model, UNI, which had a ranking score of 3.7 (ranked first in 6 tasks). To evaluate the significance of GPFM’s ranking score relative to other FMs, we performed the Nemenyi statistical test [40] (Fig. 2d). The results demonstrate that GPFM exhibited a statistically significant critical difference compared to the other eight models.
We calculated the average evaluation metrics across all 72 tasks (Fig. 2b), revealing that GPFM achieved the highest average score of 0.833, surpassing the second-best model, UNI, which scored 0.818. To assess statistical significance, we conducted a Wilcoxon signed-rank two-sided test [40] comparing GPFM with the second- and third-best models. The results showed that all p-values were below 0.001, confirming that GPFM consistently and significantly outperformed the existing FMs. Considering both the ranking perspective and the average metric aspect, the results clearly indicate that GPFM achieves state-of-the-art performance and is much more generalizable compared to the other FMs.
@@ 2.1 WSI Classification
WSI classification is pivotal in accurate cancer diagnosis. It aids in categorizing the specific subtype of cancer, which can be significantly improved by utilizing FMs. Therefore, it is important to evaluate the representation learning capabilities of different FMs. We conducted experiments on a total of 36 tasks, including 20 internal validation datasets and 16 external validation datasets. The detailed experimental results are presented in Extended Data Table 1-18.
Across 36 WSI classification tasks, ranked according to the Area Under the Curve (AUC) metric, GPFM achieved an outstanding average ranking score of 1.22, significantly surpassing the second-best model, UNI, which attained an average ranking score of 3.60 (Fig. 3a).
We assessed overall performance using average metrics: AUC, balanced accuracy, and weighted F1 score.
Specifically, GPFM achieved the highest average AUC of 0.891, a 1.6% improvement over UNI (P < 0.001; Fig. 3d).
Similarly, GPFM outperformed UNI in balanced accuracy (0.752, +3.1%, P < 0.001; Fig. 3b) and weighted F1 score (0.736, +3.0%, P < 0.001; Fig. 3c).
Additionally, GPFM achieved the best performance in both internal and external tasks, with AUCs of 0.938 (+1.6% over UNI, Fig. 3e) and 0.832 (+1.5% over UNI, Fig. 3f). These results across multiple metrics highlight GPFM’s strong generalization capability and potential for WSI classification tasks.
GPFM Enhances Diagnostic Accuracy Across Multiple Cancer Types.
GPFM demonstrates superior diagnostic accuracy across a range of cancer types and tasks. In breast cancer, GPFM outperforms other models in all six evaluated tasks, including five subtyping tasks (Fig. 3j, Extended Data Fig. 7a) and one metastasis detection task (Extended Data Fig. 7d).
For lung cancer, GPFM excels in three subtyping tasks, two metastasis detection tasks, and two primary site prediction tasks (Extended Data Fig. 7b, f, and h), with the exception of one external validation for lung cancer metastasis detection, where UNI performs slightly better.
In gastric cancer, GPFM achieves the best performance in six out of nine tasks, including vascular invasion detection (Fig. 3h), perineural invasion detection, and Lauren subtyping (Extended Data Fig. 7g and i).
Furthermore, GPFM consistently delivers top performance in tasks involving other organs, such as brain tumor subtyping, head and neck cancer primary site and T stage prediction, colon lesion grading, prostate cancer grade assessment, ovarian cancer subtyping (Extended Data Fig. 3c, d, b, c, i, and g), and renal cell carcinoma classification (Fig. 3g).
Overall, GPFM establishes itself as a leading model in cancer diagnosis across diverse tasks and cancer types.
@@
We conducted experiments on lung cancer and brain cancer slides.
GPFM achieved the best results in both TP53 mutation prediction for lung cancer, with an AUC of 0.855 (+1.3% over Phikon; Extended Data Fig. 7e), and IDH1 mutation prediction for glioma, with an internal AUC of 0.986 and an external AUC of 0.943 (Extended Data Fig. 8a).
These results, along with the cancer diagnosis findings, highlight GPFM’s superior generalizability compared to existing FMs. A key factor in this success is GPFM’s ability to integrate knowledge from expert models through a unified knowledge distillation mechanism. Unlike previous FMs that did not employ knowledge distillation, GPFM leverages this approach to learn from a broader range of data and perspectives, significantly enhancing its performance. This capability underscores GPFM’s advanced adaptability and effectiveness across diverse tasks.
@@ 2.2 Survival Analysis
Accurate prediction of a patient’s survival risk can enable more targeted and effective treatment strategies.[43, 44, 45, 46]. A robust FM is essential for improving the precision of survival risk prediction, ultimately leading to better patient outcomes. To evaluate the performance of various FMs in survival analysis, we conducted experiments on 15 datasets. Following the methodologies of previous works [47, 44, 46], we adopted the Concordance Index (C-Index) as the evaluation metric to compare the performance of different FMs.
Across the 15 survival analysis tasks, the GPFM achieved an impressive average ranking score of 2.1, ensuring the best or second-best performance in 13 tasks (Fig. 4a, d-f). In comparison, the second-best performing model, UNI, attained an average ranking score of 4.6, achieving top-2 performance in only 4 tasks (Fig. 4a, d-f). Furthermore, when evaluated using the widely recognized C-Index metric, the GPFM emerged as the top performer, achieving an average C-Index of 0.665 (Fig. 4b). This result represents a statistically significant improvement of 3.4% over UNI (P < 0.001), further demonstrating the superior generalization capability of GPFM for survival analysis tasks. To further validate the generalization of FMs, we conducted additional validation studies, including one external validation for head and neck cancer (TCGA-HNSC) and one internal validation for lung adenocarcinoma (TCGA-LUAD). In the head and neck cancer survival prediction task, UNI achieved the best performance in both the TCGA-HNSC and HANCOCK cohorts, while our method ranked as the second-best performer (Fig. 4c). However, in the lung adenocarcinoma task, GPFM demonstrated a 10.6% improvement in the CPTAC-LUAD cohort (Extended Data Fig. 9h) compared with UNI.
@@
It is noteworthy that survival analysis tasks are inherently more challenging than WSI classification, and no single model has been able to dominate these tasks (Fig. 2e). The experimental results from both WSI classification and survival analysis highlight the limited generalization capability of existing FMs. This limitation is likely attributable to the data distribution of their training sets and the pretraining methods they employ. While existing FMs exhibit limited generalization, they demonstrate exceptional performance on specific types of tasks. By leveraging their individual strengths, it is possible to construct a more powerful and versatile model. This is precisely what we have achieved in this study: we propose a unified distillation framework to distill the capabilities of existing models—particularly in tasks where they excel—into the GPFM, thereby significantly enhancing its generalization ability.
@@ 2.3 ROI Classification
The performance of WSI classification is influenced by both the feature extractor (i.e., FM) and the multiple instance learning (MIL) method. Unlike WSI classification, Region-of-Interest (ROI) classification tasks allow for a direct assessment of the FMs’ feature representation capabilities, independent of MIL methods. To this end, we employed a linear probe approach, as outlined in [48], to evaluate the FMs. Our assessment spanned 16 ROI classification tasks, encompassing 13 internal and 3 external validation datasets. Comprehensive findings from these evaluations are cataloged in Extended Data Tables 24-36.
GPFM emerged as the top performer across all 16 ROI classification tasks, securing the best ranking score of 1.88, significantly outperforming the second-ranked model, Prov-Gigapath, which scored 3.09 (Fig. 5a). In terms of conventional metrics, GPFM achieved the highest average AUC of 0.946 (+0.2% over Prov-Gigapath, P<0.001; Fig. 5d), the best weighted F1 score of 0.865 (+0.9%, P<0.001; Fig. 5c), and the highest balanced accuracy of 0.866 (+1%, P<0.001;Fig. 5b). GPFM exhibited outstanding performance in several tasks, including the detection of metastatic tissue in breast cancer (Fig. 5g), tissue type classification in lung cancer (Fig. 5h), the classification of tumor-infiltrating lymphocytes (TILs) (Fig. 5j), and the classification of gastric cancer tissues (Fig. 5k). In relatively simpler ROI classification tasks, GPFM shared the top rank with other FMs. For instance, in pancancer tissue classification (Extended Data Fig. 9f), breast tumor classification (Extended Data Fig. 9b), colorectal cancer tissue classification (Fig. 5f), and kidney tissue classification (Fig. 5e), GPFM achieved performance on par with other leading FMs. In tasks where GPFM did not achieve the top performance, it consistently ranked as the second-best method (Extended Data Fig. 9a, d, e; Fig. 5i) or the third-best method (Extended Data Fig. 9c). This consistent high ranking across diverse tasks contributed to GPFM’s overall superior performance. In addition, the average ranking scores (Fig. 5a) of UNI and Prov-Gigapath are close, with ranking scores of 3.2 and 3.1, respectively. This indicates that no single existing model dominates ROI classification tasks. In contrast, by integrating knowledge from all adopted expert models, the unified knowledge distillation enables GPFM to surpass the performance of individual models, achieving a significantly lower average ranking score of 1.88, outperforming the next-best model by more than one point. This underscores GPFM’s strength as a highly generalizable FM.
@@
Furthermore, to evaluate the robustness of GPFM in handling images with varying resolutions, we visualized the heatmap of attention scores between the [patch] tokens and [CLS] tokens of the ViT transformer (Extended Data Fig. 9g). Across four resolutions—224224, 448448, 896896, and 13441344—we observed consistent attention patterns, highlighting GPFM’s robustness in adapting to different image resolutions.